OpenStudy (anonymous):

57. What is the percentage of water in the hydrate CoCl2 • 6H2O? Show your work. (3 points)

OpenStudy (paxpolaris):

find the molecular weights of \(\large CoCl_2\) & \(\large H_2O\)

OpenStudy (unklerhaukus):

Percentage mass,

OpenStudy (paxpolaris):

\[{6(H_2O) \over (CoCl_2)+6(H_2O)} \times 100\%\]

OpenStudy (unklerhaukus):

in rough figures im getting a bit under 50%

OpenStudy (unklerhaukus):

\[(\% w_{H_2O}/w_{{CoCl}_2\cdot 6H_2O})=\frac{M_{6H_2O}}{M_{{{CoCl}_2\cdot 6H_2O}}}\]\[={6M_{H_2O} \over M_{CoCl_2}+6M_{H_2O}}\]\[=\frac{6\times(2\times M_H+M_O)}{M_{Co}+2\times M_{Cl}+6\times(2\times M_H+M_O)}\] \[M_H=1.01\text u\]\[M_O=16.00 \text u\]\[M_{Co}=58.93 \text u\]\[M_{Cl}=35.45\text u\]